# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### v8

@@ 3.4 Data Description
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These are used to test the smoothness of the consumptions in each subcategory. To test the roles of risk-coping strategies, I consider five net income categories: savings, insurance, borrowing, credit purchase, and gifts. Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings. I also consider the sales of miscellaneous goods and labor earnings from the family members except for the head to test the contribution of the asset sales and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^47] Online Appendix B.9 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^48] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping behaviors (e.g., Udry 1994, 1995). Since this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^49] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households had been stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts had rarely occurred among the THBS households. Although family size is included in all regressions to be conservative, it has little effect on the results. [^50] The family size variable shall be omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations. Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and are absorbed by the household fixed effect. Panel C of Table 3 shows the summary statistics on the family size variable.
@@
(the other family members)
@@
Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33-and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of assets” includes the sale of daily miscellaneous goods such as newspapers and empty bottles.
Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar to those for the semi-monthly panels.
Panel D: The summary statistics of the number of days with illness for family members among the 33-panel units in the semi-monthly panel dataset are listed. Since both are lagged variables, the initial time period in all the units is omitted. Similarly, a household with two semi-month observations is excluded. Similarly, .
Sources: Created by the author from the THBS sample.
@@
Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure. Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in the nominal income during this period. This trend in earnings was, however, offset by the steep inflation under the expansions after the war (Section 2.1). In Figure 4d, the deviation of expenditure from income in January must capture the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The humps in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans. These support the evidence that the budget data from the THBS households can capture a large part of the macroeconomic trends.
@@
Online Appendix Figure B.11 illustrates the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. [^51] Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient. I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks because the daily panels include a set of consecutive cells with no income. While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers had experienced idiosyncratic income shocks in some cells. Regardless of favorable or adverse shocks, these cells that deviated from mean tendency are helpful in investigating how those households coped with the idiosyncratic shocks. Despite this, it is helpful to focus on the illness as it is the least predictable shock (Gertler and Gruber 2002). Although joblessness might be another shock, the THBS heads rarely changed their occupation during the survey period as they are the skilled workers who were less likely to move to other jobs. [^52] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey. However, I confirmed that the households with longer and shorter time-series observations have similar family characteristics (Online Appendix B.3). Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk. Moreover, the number of dismissals at the Tsukishima factories seems to be small. In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory for which statistics are available from the official report of the Tsukishima Survey was only 1.5% (Department of Health, Ministry of the Interior 1923a, p. 415). In contrast, fortunately, the illness among the THBS family is documented systematically. As explained in Section 2.3, neither the comprehensive health insurance system nor the factory’s compensation for sickness was available. Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates. [^53] See Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan. The official report mentioned that the workers in the machinery factories in Tsukishima had faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p. 14).
Considering this, I first provide the overall results from the specification using the changes in earnings as the idiosyncratic income shock in Sections 4.1 and 4.2. I then consider the structure behind the adverse income shocks using the heads’ temporary illness to investigate the households’ consumption and risk-coping behaviors in Section 4.3. Panel D of Table 3 shows the summary statistics on the health shock variables.
@@
This study investigates consumption smoothing behaviors among the skilled factory workers’ households in the machinery sector in Tsukishima. This empirical setting is preferable because the different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks in the spirit of Schulhofer-Wohl (2011). Specifically, I consider the case that the heads who worked in large-scale or governmental factories to have potentially different risk preferences because those workers might have lower risk tolerance than those who worked in smaller enterprises. I confirmed that the sensitivities to the aggregate shocks are materially similar in both groups in the statistical sense. [^54] Online Appendix B.10 summarizes the finer details of this analysis. In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and government factories, family size variable, and household fixed effects. The results are unchanged if I trim a few semi-months with a smaller number of cross-sectional observations. This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery sectors had different risk preferences. [^55] The skilled workers in the smithing sector might be a business owner (Online Appendix A.6). Similarly, the workers in the non-machinery sectors in Tsukishima, a representative machinery production area, might have had a different, presumably higher, risk-tolerance. This supports the evidence that conditional on the household fixed effects, it is plausible to assume that the THBS households shared similar risk preferences.
@@ 4.1 Consumption Smoothing
I consider a linear fixed-effects model designed to test the consumption smoothing. [^56] Another empirical specification used to test the consumption smoothing is the first-difference model, including the aggregate consumption to capture the macroeconomic shocks in the economy (Mace 1991). In other words, this alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. Although the consumption of the THBS households tends to capture the macroeconomic trend in the consumption (Section 3.3), those households do not cover entire households in the target economy, Tsukishima. This disturbs the interpretation of the coefficient on aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for the macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ shows income elasticity, ranging from zero for perfect insurance to one for the absence of insurance. [^57] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption smoothing literature. Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences setting (TWFE-DID) yet. In fact, the bias-corrected estimator of the TWFE-DID with continuous treatment for multiple periods is underdeveloped yet (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review). The cluster-robust variance-covariance matrix estimator is used to deal with heteroskedasticity and serial dependency (Arellano 1987).
@@ Results
Notes: This table shows the results of equation 1: the regressions of the 11 measures of log-transformed consumption on log-transformed disposable income as well as on the family size control, household-fixed effect, and time-fixed effect. Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.358$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumption. An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant. This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to deal with those shocks to smooth food consumption. The estimates for the entertainment and transportation expenses ($0.307$ and $0.278$) are also highly statistically significant. This may make sense given that those expenses include family travel and theatergoing. [^58] The workers can enter their factories by walking in Tsukishima. Thus, the costs of commuting to the factory are not usually documented in the budget books. Similarly, the estimate for gifts shows a large magnitude ($0.523$), reflecting that the expenses on the others are essentially optional. The results for the other subcategories are statistically insignificant or unclear. The estimates for furniture and clothes are weakly statistically significant. Although less than conclusive, this may reflect that both categories are luxury categories. The estimate for the education category is small but also weakly statistically significant ($0.101$). This may be because it includes the allowance for children, which could be easily reduced. [^59] Toyotaro Miyoshi served as a research assistant at the Tsukishima Survey while he was a student at the Tokyo Imperial University. He left important retrospectives on the survey in his works (Online Appendix B.1). According to his book, since there were no nurseries, children in these working-class households regularly received a small allowance to play outside while their parents were working (Miyoshi 1989, pp. 383–384; 386; 387).
@@
I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed in a longer span. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption is now $0.433$, which is roughly $1.2$ times greater than that from the semi-monthly series. One potential explanation behind this difference is that the available sources of temporary income have been limited, and those sources are not useful for compensating for adverse income shocks lasting for at least one month. The most noticeable example is the food category, which comprises the largest portion of the expenditure (Panel A of Table 3). The estimate is $0.293$, more than two times greater than the estimate for the semi-monthly series ($0.136$). This implies that the households’ food expenditures are likely to be insured in a few weeks but not over one month. As I will delve into the mechanism behind this behavior in the following subsections, credit purchases may work to smooth food consumption in the short run. Similarly, the expenses for entertainment and transportation are much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series. The estimate for gifts is slightly smaller but still in a similar range. In short, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets. The expenses on those unnecessary categories might particularly be easily cut in response to relatively persistent shocks.
@@
In contrast, a few subcategories are not sensitive to the idiosyncratic shocks over different time-series frequencies. For example, the estimates for the housing, utilities, and medical categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect that the rent and utility payments were relatively fixed. [^60] For example, rent payment is usually fixed and paid once per month. Online Appendix C.2 shows that the housing category should be analyzed based on the adjusted monthly series rather than the semi-monthly series. The medical expenses included the fee for bathing (yuya). People had taken both every few days at that time, which was less likely to be cut, albeit hairdressing for women (kamiyui) in the same category may be likely to be cut. The estimates for the furniture, clothes, and education categories are statistically insignificant in the adjusted monthly series. While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, like unanticipated losses of household goods, clothes, and stationaries including writing papers. [^61] An alternative explanation for the clothes category relates to seasonality. The workers had bought their work clothes for the new season (for example, the OISR, Archives of the Tsukishima Survey (THBS #40, 31st March 1919)). Since those clothes are rather necessary items, this sort of seasonal expense may attenuate the estimate under low frequencey series. This systematically attenuates the estimates for these specific categories in the adjusted monthly series. [^62] Generally, the panels with short-run time bins are more likely to be influenced by unexpected shocks in consumption, which tends to attenuate the income elasticity (Nelson 1994). If this mechanism worked, subcategories that are sensitive (insensitive) to unexpected shocks in consumption should be more (less) attenuated. In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels across a large part of subcategories. Thus, on average, my result indicates that households might face more difficulties in smoothing their consumption in the longer run.
@@
Finally, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. Despite this, several subcategories take values close to one or large negative values, which are no longer economically interpretable. [^63] For example, the estimate for housing is now negative. The households paid a fixed rent amount monthly. Given that the rent payday is fixed, for example, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these particular months may have created a strong but false positive correlation. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives in most of the consumption categories and may cause critical issues in statistical inference for a few subcategories.

### v10

@@ 3.4 Data Description
The THBS budget book has two primary sections: expenditure and income. I digitized records in each category following the categorization used in the official report. Panel A of Table 3 lists the $11$ subcategories for expenditures: food, housing, utilities, furniture, clothes, education, medical, entertainment, transportation, gifts, and miscellaneous. These are used to test the smoothness of the consumptions in each subcategory. To test the roles of risk-coping strategies, I considered five net income categories—savings, insurance, borrowing, credit purchase, and gifts. Net variables are defined as the income minus expenditure in each category; for example, net savings refers to the difference between the withdrawal amount and deposits to savings. I also considered the sales of miscellaneous goods and labor earnings from the family members (except for the head) to test the contribution of the sales of small assets and intrahousehold labor-supply adjustments. Panel B of Table 3 summarizes these variables. [^43] Online Appendix B.9 presents the results for the panel unit root tests. The null hypothesis of unit roots in all the panels is rejected at the conventional level for all these variables.
Alterations in family size may shift the household’s preferences (Jappelli and Pistaferri 2017). [^44] In village economies, landholding and asset stocks such as livestock and grain may also be correlated with the household’s risk-coping strategies (e.g., Udry 1994, 1995). As this study analyzes urban factory workers’ households, the roles of these assets can be discarded. Fortunately, the archives of the Tsukishima Survey include a set of sheets documenting the changes in the composition of the THBS households (Kazoku idō hyō). [^45] The Ohara Institute for Social Research, Archives of the Tsukishima Survey (THBS, unreleased). I found that the size of households was stable over time: one household experienced changes in the number of family members due to the birth of a son during the sample periods. This supports the evidence that the preference shifts rarely occurred among the THBS households. Although family size was included in all regressions, to be conservative, it had little effect on the results. [^46] The family size variable was omitted in several regressions due to the multicollinearity in the fixed-effects models given its small within variations. Although the number of family members in different age bins is available from the THBS, those family composition variables are also time-constant and were absorbed by the household fixed effect. Panel C of Table 3 shows the summary statistics on the family size variable.
@@
(other family members)
@@
Panel B: This panel shows the summary statistics of the variables used to test the risk-coping strategies for the 33- and 26-panel units in the semi-monthly and adjusted monthly panel datasets, respectively. Each net income variable is defined as the difference between income and expenses. “Net savings” refers to the difference between the amount withdrawn and the total deposits. “Net insurance” refers to the amount received from insurance minus the expenses paid for insurance. “Net borrowing” refers to the amount of borrowing minus the debt payments. “Net credit purchase” refers to the amount of credit purchases minus credit redeemed. “Net gifts” refers to the total received amounts, including both pecuniary and non-pecuniary gifts, minus the payments for gifts. “Other family member’s earnings” refers to the total income earned by all family members except for the head of the household. “Sales of miscellaneous assets” includes the sale of miscellaneous goods such as old newspapers and empty bottles.
Panel C: The summary statistics of the family size variable for the 33-panel units in the semi-monthly panel dataset are listed. The statistics for the adjusted monthly panel dataset are not reported because they are materially similar.
Panel D: The summary statistics of the number of days with illness for family members among the 33-panel units in the semi-monthly panel dataset are listed. As those are lagged variables, the initial period is omitted in all the units. Accordingly, a household with two semi-month observations is excluded.
Source: Created by the author from the THBS sample.
@@
Next, I illustrate the time-series figures of the THBS dataset to overview the trend in income and expenditure. Figure 4c shows a weak increasing trend in the average income, which is consistent with the overall macroeconomic trend in nominal income during this period. This trend in earnings was, however, offset by steep inflation caused by expansions after the war (Section 2.1). In Figure 4d, the deviation of expenditure from income in January captures the expenses on the goods for the New Year events. The gaps in June to the first half of July may reflect the payments of bonuses. The peaks in March may reflect the preparation for the new fiscal and academic year from April, suggesting a higher proportion of temporary income from withdrawals and loans. These show that the budget data from the THBS households can capture a substantial aspect of the macroeconomic trends.
@@
Online Appendix Figure B.11 illustrates the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. [^47] Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as $\Delta\hat{\upsilon}_{i,t}=\Delta\textit{Income}_{i,t}-\hat{\alpha}\Delta\overline{\textit{Income}}_{.,t}$, where $\overline{\textit{Income}}_{.,t}=\sum_{i=1}^{n_{t}}\frac{\textit{Income}_{i,t}}{n_{t}}$ and $\hat{\alpha}$ is the estimated coefficient. While the residuals satisfy the zero-average property, they occasionally take large negative and positive values, suggesting that the workers experienced idiosyncratic income shocks in some cells. Regardless of favorable or adverse shocks, these cells that deviated from the mean tendency help investigate how those households coped with the idiosyncratic shocks. Despite this, it is useful to consider illness as it is the least predictable shock (Gertler and Gruber 2002). Fortunately, the illness among the THBS families is documented systematically and thus, available for analysis. As explained in Section 2.3, neither a comprehensive health insurance system nor compensation for illness from the factories was available. Moreover, the sanitary conditions in the cities were low, leading to higher mortality rates. [^48] See Schneider and Ogasawara (2018) for the overviews on the sanitary conditions in interwar Japan. The official report mentioned that the workers in the machinery factories in Tsukishima faced the risk of diseases under severe sanitary conditions (Department of Health, Ministry of the Interior 1923a, p. 14). Although unemployment might be another shock, the THBS heads rarely changed their occupation during the survey period as they were skilled workers who were less likely to move to other jobs. [^49] Another possibility is that the THBS heads who changed their jobs might have been omitted from the survey. However, I confirmed that the households with longer and shorter time-series observations had similar family characteristics (Online Appendix B.3). Thus, this omission does not lead to selection bias in terms of the potential vulnerabilities to the idiosyncratic risk. Moreover, the number of dismissals at the Tsukishima factories would be small. In fact, the average monthly dismissal rate (i.e., the number of dismissals/workers) from June 1919 to May 1920 at the one factory available from the official report of the Tsukishima Survey was only $1.5$% (Department of Health, Ministry of the Interior 1923a, p. 415). Panel D of Table 3 shows the summary statistics on the health shock variables.
@@
The different attitudes toward risks among households with different occupations and regions may lead to bias in the standard risk-sharing regressions (Schulhofer-Wohl 2011; Mazzocco and Saini 2012). In contrast, this study investigates consumption-smoothing strategies among homogeneous households — skilled factory workers’ households in the machinery manufacturing sector in Tsukishima. To assess the homogeneity of the risk preference among THBS households, I tested whether the earnings of the heads who worked in a few specific factories or sectors showed different responses to the aggregate shocks, following Schulhofer-Wohl (2011). Specifically, I consider the case that the heads who worked in large-scale or governmental factories might have had different risk preferences because those workers might have had lower risk tolerance than those who worked in smaller enterprises. I confirmed that the sensitivities to the aggregate shocks are statistically similar. [^50] Online Appendix B.10 summarizes the finer details of the analysis. In short, the heads’ earnings are regressed on the aggregate consumption, its interaction term with respect to the indicator variable for the large-scale and governmental factories, family size variable, and household fixed effects. The results are unchanged if I trim a few semi-months with a smaller number of cross-sectional observations. This result is unchanged if I assume that the heads who worked in the smithing factory or in the non-machinery manufacturing sectors had different risk preferences. [^51] The skilled workers in the smithing sector could be business owners (Online Appendix A.6). Similarly, the workers in the non-machinery manufacturing sectors in Tsukishima—a representative machinery production area—could have a different, presumably higher, risk tolerance. This shows that it is plausible to assume that the THBS households shared similar risk preferences, conditional on the household fixed effects.
@@ 4.1 Consumption Smoothing
I consider a linear fixed-effects model to test consumption smoothing. [^52] Another empirical specification used to test consumption smoothing is the first-difference model, including the aggregate consumption to capture macroeconomic shocks in the economy (Mace 1991). This alternative model requires the interpretation of the estimated coefficient on the aggregate measure of consumption. However, the THBS households do not include entire households in the target economy—Tsukishima. This may disturb the interpretation of the coefficient on aggregate consumption. To be conservative, therefore, I use the two-way fixed-effects model to control for macroeconomic shocks (Cochrane 1991; Ravallion and Chaudhuri 1997). Online Appendix C.1 provides the conceptual framework and the derivation of my empirical specification. For household $i$ in time $t$, the reduced form equation is characterized as follows:
@@
where $c_{i,t}$ is consumption, $y_{i,t}$ is disposable income, $x_{i,t}$ is the family size control, $\mu_{i}$ is the household fixed effect, $\phi_{t}$ is the semi-month fixed effect, and $u_{i,t}$ is a random error term. The household fixed effect captures the time-constant unobservable factors such as permanent income and consumption preference. The time-fixed effect effectively removes the unobservable macroeconomic shocks and trends described in Section 3.4. The estimate of $\gamma$ shows income elasticity, ranging from zero for perfect insurance to one for the absence of insurance. [^53] This is comparable to the estimate from the same conceptual framework: this comparability is important in the consumption-smoothing literature. Note that relevant previous studies have not integrated the recent theoretical discussion on the fixed effect model with two-way error components under the difference-in-differences (TWFE-DID) setting yet. In fact, the bias-corrected estimator for the TWFE-DID model with continuous treatment for multiple periods is underdeveloped (see de Chaisemartin and D’Haultfœuille (2023) for a comprehensive review). The cluster-robust variance-covariance matrix estimator is used to manage heteroskedasticity and serial dependency (Arellano 1987).
@@ Results
Notes: This table shows the results of Equation 1. The regressions of the 11 measures of log-transformed consumption on log-transformed disposable income, the family size control, household-fixed effect, and time-fixed effect are listed. Columns 1, 2, and 3 show the results for the semi-monthly, adjusted monthly, and unadjusted monthly panel datasets, respectively. The estimated coefficients on log-transformed disposable income are listed in the columns named “Coef.”.
@@
Column (1) of Table 4 presents the results from the semi-monthly panel dataset. In the total consumption, the estimated coefficient is $0.358$. This means that a one percentage point decrease in disposable income decreases the total expenditure by $0.36$ percentage point, suggesting that the factory worker households could not thoroughly smooth their consumption. An important finding is that the estimate for food categories is smaller ($0.136$) but highly statistically significant. This implies that while the households tended to reduce their food expenses when they faced adverse income shocks, they might have had some means to manage those shocks to smooth food consumption. The estimates for the entertainment and transportation expenses ($0.307$ and $0.278$) are also highly statistically significant. This may be understandable because those expenses include family travel and theatergoing. [^54] The workers could walk to the factories in Tsukishima, so the costs of commuting were not usually documented in the budget books. Similarly, the estimate for gifts shows a large amount ($0.523$), reflecting that the expenses on the others were optional. The results for the other subcategories are statistically insignificant or unclear. The estimates for furniture and clothes are weakly statistically significant. Although not conclusive, this may reflect that both categories were luxury categories. The estimate for the education category is small but also weakly statistically significant ($0.101$). This may be because it includes an allowance for children, which could be easily decreased. [^55] Toyotaro Miyoshi served as a research assistant at the Tsukishima Survey while he was a student at the Tokyo Imperial University. He left important retrospectives on the survey in his works (Online Appendix B.1). According to his book, since there were no nurseries, children in these working-class households regularly received a small allowance to play outside while their parents were working (Miyoshi 1989, pp. 383–384; 386; 387).
@@
I present the results for the adjusted monthly panel dataset in Column (2) to see whether the income elasticities changed over a longer span. Not surprisingly, the adjusted monthly series yields greater estimates than the semi-monthly series. The estimate for total consumption was $0.433$, which is approximately $1.2$ times greater than that from the semi-monthly series. One potential explanation for this difference is that the available sources of temporary income might have been limited, and those sources were not useful for compensating for adverse income shocks that lasted for at least one month. The most noticeable example is the food category, which comprised the largest portion of the expenditure (Panel A of Table 3). The estimate is $0.293$, more than twice the estimate for the semi-monthly series ($0.136$). This implies that the households’ food expenditures were likely to be insured for a few weeks but not over one month. I will investigate the mechanism underlying this behavior in the following subsections, and credit purchases may be used to smooth food consumption in the short run. Similarly, the expenses for entertainment and transportation were much more sensitive ($0.716$ and $0.605$) to the shocks in the adjusted monthly series. The estimate for gifts is slightly smaller but in a similar range. In short, while the household could compensate for their income losses for a few weeks, they could not do so for several weeks, presumably because of borrowing constraints and/or a limited amount of liquidity assets. The expenses on those unnecessary categories might particularly be cut in response to relatively persistent shocks.
@@
In contrast, a few subcategories were not sensitive to the idiosyncratic shocks over different time-series frequencies. For example, the estimates for the housing, utilities, and medical categories are not statistically significant in both semi-monthly and adjusted monthly series. This may reflect the fact that the rent and utility payments were relatively fixed. [^56] For example, rent payment was usually fixed and paid once per month. Online Appendix C.2 shows that the housing category should be analyzed based on the adjusted monthly series rather than the semi-monthly series. The medical expenses included the fee for bathing (yuya). People took baths every few days at that time, which was less likely to be cut, although hairdressing for women (kamiyui) in the same category was likely to be cut. The estimates for the furniture, clothes, and education categories are statistically insignificant in the adjusted monthly series. While this could partly be the efficiency issue under a smaller sample size, those expenses may be sensitive to unexpected shocks in consumption, such as unanticipated losses of household goods, clothes, and stationery, including writing paper. [^57] An alternative explanation for the clothes category relates to seasonality. The workers had bought their work clothes for the new season (e.g., the OISR, Archives of the Tsukishima Survey (THBS #40, March 31, 1919)). As those clothes were necessary items, those seasonal expenses may attenuate the estimate. This may systematically attenuate the estimates for the specific categories in the adjusted monthly series. [^58] Generally, the panels with short-run time bins were more likely to be influenced by unexpected consumption shocks, which attenuated income elasticity (Nelson 1994). If this mechanism worked, subcategories that were sensitive (insensitive) to unexpected consumption shocks should be more (less) attenuated. In my result, however, the estimates for the adjusted monthly panels are greater than those for the semi-monthly panels in a large part of subcategories, suggesting that households faced more difficulties with smoothing their consumption in the long run.
@@
Finally, I assess the validity of the aggregation method used in this study. Adjusting the aggregate data to account for the timing of paydays reveals more transparent relationships between income and expenditure than the unadjusted data (Section 3.3). To demonstrate this mechanism, I listed the estimates from the unadjusted monthly panel dataset in Column (3). The estimate for total consumption is $0.26$, which is much lower than the estimate for the adjusted series ($0.43$). In addition, most subcategories show systematically smaller estimates than those from the adjusted data listed in Column (2). For example, the estimate for the food category is now $0.10$ and statistically insignificant. This means that measurement errors in the unadjusted aggregation lead to strong attenuation bias, which results in false negatives (i.e., false rejections of the null of full risk-sharing hypothesis) in most of the consumption categories. I, therefore, specified a generating process of systematic measurement errors in the aggregate survey data, which has attracted attention in the macroeconomics literature (Cochrane 1991; Nelson 1994; Dynarski et al. 1997; Gervais and Klein 2010). Moreover, several subcategories select values close to one or large negative values, which are no longer economically interpretable. [^59] For example, the estimate for housing was negative. The households paid a fixed rent amount monthly. Given that the rent payday was fixed, the shifts in wage paydays might create cells with less income with a fixed amount of rent. In the case of seasonal consumption categories, the shifts may coincide with the timing of seasonal spending. The THBS households tended to purchase work clothes in March to prepare for the start of the new fiscal year in April. They also spent more on clothing in June, likely to buy summer clothes. The shift of paydays to these months may have created a strong but false positive correlation. This means that the miss-assignment of paydays in the unadjusted data may cause critical statistical inference issues in a few subcategories.